# ODIN: A Single Model for 2D and 3D Perception

### v1

@@ Abstract
State-of-the-art models on contemporary 3D perception benchmarks like ScanNet consume and label dataset-provided 3D point clouds, obtained through post processing of sensed multiview RGB-D images. They are typically trained in-domain, forego large-scale 2D pre-training and outperform alternatives that featurize the posed RGB-D multiview images instead. The gap in performance between methods that consume posed images versus post-processed 3D point clouds has fueled the belief that 2D and 3D perception require distinct model architectures. In this paper, we challenge this view and propose ODIN (Omni-Dimensional INstance segmentation), a model that can segment and label both 2D RGB images and 3D point clouds, using a transformer architecture that alternates between 2D within-view and 3D cross-view information fusion. Our model differentiates 2D and 3D feature operations through the positional encodings of the tokens involved, which capture pixel coordinates for 2D patch tokens and 3D coordinates for 3D feature tokens. ODIN achieves state-of-the-art performance on ScanNet200, Matterport3D and AI2THOR 3D instance segmentation benchmarks, and competitive performance on ScanNet, S3DIS and COCO. It outperforms all previous works by a wide margin when the sensed 3D point cloud is used in place of the point cloud sampled from 3D mesh. When used as the 3D perception engine in an instructable embodied agent architecture, it sets a new state-of-the-art on the TEACh action-from-dialogue benchmark. Our code and checkpoints can be found at the project website https://odin-seg.github.io.
@@ 1 Introduction
We observe that part of the issue lies in a key implementation detail underlying these 3D benchmark evaluations. ScanNet and ScanNet200 do not actually ask methods to use RGB-D images as input, even though this is the sensor data. Instead, these benchmarks first register all RGB-D frames into a single colored point cloud and reconstruct the scene as cleanly as possible, relying on manually tuned stages for bundle adjustment, outlier rejection and meshing, and ask models to label the output reconstruction. While it is certainly viable to scan and reconstruct a room before labelling any of the objects inside, this pipeline is perhaps inconsistent with the goals of embodied vision (and typical 2D vision), which involves dealing with actual sensor data and accounting for missing or partial observations. We therefore hypothesize that method rankings will change, and the impact of 2D pre-training will become evident, if we force the 3D models to take posed RGB-D frames as input rather than pre-computed reconstructions. Our revised evaluation setting also opens the door to new methods, which can train and perform inference in either single-view or multi-view settings, with either RGB or RGB-D sensors.
@@
We propose Omni-Dimensional INstance segmentation (${}^{\dagger}$footnotetext: ${}^{\dagger}$he Norse god Odin sacrificed one of his eyes for wisdom, trading one mode of perception for a more important one. Our approach sacrifices perception on post-processed meshes for perception on raw sensor data. , a model for 2D and 3D object segmentation and labelling that can parse single-view RGB images and/or multiview posed RGB-D images. As shown in Fig. 1, ODIN alternates between 2D and 3D stages in its architecture, fusing information in 2D within each image view, and in 3D across posed image views. At each 2D-to-3D transition, it unprojects 2D tokens to their 3D locations using the depth maps and camera parameters, and at each 3D-to-2D transition, it projects 3D tokens back to their image locations. Our model differentiates between 2D and 3D features through the positional encodings of the tokens involved, which capture pixel coordinates for 2D patch tokens and 3D coordinates for 3D feature tokens. When dealing with 2D single-view input, our architecture simply skips the 3D layers and makes a forward pass with 2D layers alone.
@@
We test ODIN in 2D and 3D instance segmentation and 3D semantic segmentation on the 2D COCO object segmentation benchmark and the 3D benchmarks of ScanNet [6], ScanNet200 [41], Matterport3D [2], S3DIS [1] and AI2THOR [25, 7]. When compared to methods using pre-computed mesh point cloud as input, our approach performs slightly worse than state-of-the-art on ScanNet and S3DIS, but better on ScanNet200 and Matterport3D. When using real sensor data as input for all methods, our method performs even better, outperforming all prior work by a wide margin, in all datasets. We demonstrate that our model’s ability to jointly train on 3D and 2D datasets results in performance increase on 3D benchmarks, and also yields competitive segmentation accuracy on the 2D COCO benchmark. Our ablations show that interleaving 2D and 3D fusion operations outperforms designs where we first process in 2D and then move to 3D, or simply paint 3D points with 2D features. Stepping toward our broader goal of embodied vision, we also deploy ODIN as the 3D object segmentor of a SOTA embodied agent model [42] on the simulation benchmark TEACh [36] in the setup with access to RGB-D and pose information from the simulator, and demonstrate that our model sets a new state-of-the-art. We make our code publicly available at https://odin-seg.github.io.
@@ 2 Related Work
Early methods in 3D instance segmentation [22, 15, 3, 30, 49, 58] group their semantic segmentation outputs into individual instances. Recently, Mask2Former [4] achieved state-of-the-art in 2D instance segmentation by instantiating object queries, each directly predicting an instance segmentation mask by doing dot-product with the feature map of the input image. Inspired by it, Mask3D [44] abandons the grouping strategy of prior 3D models to use the simple decoder head of Mask2Former. MAFT [28] and QueryFormer [34] improve over Mask3D by incorporating better query initialization strategies and/or relative positional embeddings. While this shift to Mask2Former-like architecture brought the 3D instance segmentation architectures closer to their 2D counterparts, the inputs and backbones remain very different: 2D models use popular and often pre-trained backbones [16, 33], while 3D methods [44] operate over point clouds and use sparse convolution-based backbones [5], trained from scratch on small-scale 3D datasets. In this work, we propose to directly use RGB-D input and design architectures that can leverage strong 2D backbones to achieve strong performance on 3D benchmarks.
@@
Recent advancements in 2D foundation models [39, 24] have spurred efforts to apply them to 3D tasks such as point cloud classification [38, 56, 52], zero-shot 3D semantic segmentation [37, 21, 14] and more recently, zero-shot instance segmentation [47]. Commonly, these methods leverage 2D foundation models to featurize RGB images, project 3D point clouds onto these images, employ occlusion reasoning using depth and integrate features from all views through simple techniques like mean-pooling. Notably, these approaches predominantly focus on semantic segmentation, emphasizing pixel-wise labeling, rather than instance labeling, which necessitates cross-view reasoning to associate the same object instance across multiple views. OpenMask3D [47] is the only method that we are aware of that attempts 3D instance segmentation using 2D foundation models, by simply training a class-agnostic 3D object segmentor on 3D point clouds and labelling it utilizing CLIP 2D features. Despite their effectiveness in a zero-shot setting, they generally lag behind SOTA 3D supervised methods by 15-20%. Rather than relying on features from foundation models, certain works [12, 10] create 3D pseudo-labels using pre-trained 2D models. Another line of work involves fitting Neural-Radiance Fields (NeRFs), incorporating features from CLIP [23, 48] or per-view instance segmentations from state-of-the-art 2D segmentors [46]. Despite their impressive results, these approaches require expensive per-scene optimization that prohibits testing on all test scenes to compare against SOTA 3D discriminative models. Instead of repurposing 2D foundation models for 3D tasks, Omnivore [13] proposes to build a unified architecture that can handle multiple visual modalities like images, videos and single-view RGB-D image but they only show results for classification tasks. We similarly propose a single unified model capable of performing both single-view 2D and multi-view 3D instance and semantic segmentation tasks while utilizing pre-trained weights for the majority of our architecture.
@@ 3 Method
$q_{pos}$ $=MLP(0);$ (1)
$k_{pos}=v_{pos}$ $=MLP(p_{i}-p_{j}),$ (2)
@@
Open vocabulary class decoder: Drawing inspiration from prior open-vocabulary detection methods [19, 29, 61], we introduce an alternative classification head capable of handling an arbitrary number of semantic classes. This modification is essential for joint training on multiple datasets. Similar to BUTD-DETR [19] and GLIP [29], we supply the model with a detection prompt formed by concatenating object categories into a sentence (e.g., “Chair. Table. Sofa.”) and encode it using RoBERTa [32]. In the query-refinement block, queries additionally attend to these text tokens before attending to the upsampled feature maps. For semantic class prediction, we first perform a dot-product operation between queries and language tokens, generating one logit per token in the detection prompt. The logits corresponding to prompt tokens for a specific object class are then averaged to derive per-class logits. This can handle multi-word noun phrases such as “shower curtain”, where we average the logits corresponding to “shower” and “curtain”. The segmentation masks are predicted by a pixel-/point-wise dot-product, in the same fashion as described earlier.
@@ 4.1 Evaluation on 3D benchmarks
Table 1: Evaluation on Established 3D Benchmarks. (${}^{\mathsection}$ = trained by us using official codebase)
(a) Comparison on ScanNet for Instance Segmentation Task.
@@
Sensor RGBD Point Cloud ${}^{\mathsection}$ [44] 43.9 60.0 69.9
ODIN-ResNet50 (Ours) 45.7 66.3 81.8
ODIN-Swin-B (Ours) 48.0 69.2 83.5
@@
(b) Comparison on ScanNet for Semantic Segmentation Task.
@@
ODIN-ResNet50 (Ours) 73.2
ODIN-Swin-B (Ours) 76.0
@@
(c) Comparison on ScanNet200 for Instance Segmentation Task.
@@
Sensor RGBD Point Cloud Mask3D [44] ${}^{\mathsection}$ 15.5 21.4 24.3
ODIN-ResNet50 (Ours) 26.0 37.6 43.8
ODIN-Swin-B (Ours) 30.0 43.0 51.0
@@
(d) Comparison on ScanNet200 for Semantic Segmentation Task.
@@
Sensor RGBD Point Cloud ODIN-ResNet50 (Ours) 33.9
ODIN-Swin-B (Ours) 38.2
@@
Datasets: First, we test our model on 3D instance and semantic segmentation in the ScanNet [6] and ScanNet200 [41] benchmarks. The objective in these benchmarks is to label the point cloud sampled from the 3D mesh of a scene reconstructed from raw sensor data. ScanNet evaluates on 20 common semantic classes, while ScanNet200 uses 200 classes, which is more representative of the long-tailed object distribution encountered in the real world. We report results on the official validation split of these datasets.
@@
ODIN sets a new SOTA in sensor-based semantic segmentation on ScanNet (Tab. 1(b)) outperforming all methods operating over sensor RGB point cloud including the SOTA DeepViewAgg [40]. Additionally, our model closely matches the performance of models operating on rendered RGB-D point cloud and mesh sampled point cloud.
ODIN sets a new instance segmentation SOTA on the long-tailed ScanNet200 dataset (Tab. 1(c)) outperforming SOTA 3D models on all setups including the models trained on mesh-sampled point cloud especially by a large margin in mAP25 metric, while exclusively utilizing sensor RGB-D data. This highlights the contribution of 2D features, particularly in detecting a long tail of class distribution where limited 3D data is available. We show more detailed results with performance on the head, common and tail classes in the appendix ( Sec. A.3).
@@
Table 2: Comparison on AI2THOR for Semantic and Instance Segmentation.
@@
ODIN-ResNet50 (Ours) 64.9 74.1 80.0 71.2
ODIN-Swin-B (Ours) 67.6 76.5 81.8 74.2
@@ 4.2 Evaluation on multiview RGB-D in simulation
Using the AI2THOR [25] simulation environment with procedural homes from ProcThor [7], we collected RGB-D data for 1500 scenes (1200 training, 300 test) of similar size as ScanNet (more details in appendix,  Appendix B). We train and evaluate our model and SOTA Mask3D [44] on the unprojected RGB-D images. As shown in Tab. 2, our model outperforms Mask3D by 7% mAP, showing strong performance in a directly comparable RGB-D setup. It suggests that current real-world benchmarks may restrain models that featurizes RGB-D sensor point clouds due to misalignments. We hope this encourages the community to also focus on directly collecting, labeling, and benchmarking RGB-D sensor data.
@@ 4.3 Embodied Instruction Following
Mask3D [44]
43.9 60.0 69.9
Mask2Former [4]
43.7
ODIN (trained in 2D)
43.6
ODIN (trained in 3D)
45.7 66.3 81.8
ODIN (trained jointly)
48.3 69.5 82.2 40.7
@@
Model
mAP mIoU
ODIN (Ours)
45.7 73.2
No 3D Fusion
38.8 71.9
No interleaving
40.0 72.5
@@
Model
mAP mIoU
ODIN (Ours)
45.7 73.2
Only pre-trained backbone
41.0 71.6
No pre-trained features
36.6 65.4
@@
Model
ResNet50 Swin-B
@@
ODIN (Ours)
45.7 73.2 48.0 76.0
With frozen backbone
45.7 72.6 45.6 74.7
We apply ODIN in the embodied setups of TEACh [36] and ALFRED [45] where agents have access to RGB, depth and camera poses and need to infer and execute task and action plans from dialogue segments and instructions, respectively. These agents operate in dynamic home environments and cannot afford expensive mesh building steps. Detecting objects well is critical for task success in both cases. Prior SOTA methods [42, 36] run per-view 2D instance segmentation models [4, 9] and link the detected instances using simple temporal reasoning regarding spatial and appearance proximity. Instead, ODIN processes its last $N$ egocentric views and segments objects instances directly in 3D. We equip HELPER [42], a state-of-the-art embodied model, with ODIN as its 3D object detection engine. We evaluate using Task Sucess Rate (SR) which checks if the entire task is executed successfully, and Goal Conditioned Success Rate (GC) which checks the proportion of satisfied subgoals across all episodes [45, 36]. We perform evaluation on ”valid-seen” (houses similar to the training set) and ”valid-unseen” (different houses) splits. In Tab. 3, we observe that HELPER with ODIN as its 3D object detector significantly outperforms HELPER that uses the original 2D detection plus linking perception pipeline.
Figure 3: Qualitative Results of ODIN on 3D and 2D inputs.

### v2

# ODIN: A Single Model for 2D and 3D Segmentation

@@ Abstract
State-of-the-art models on contemporary 3D segmentation benchmarks like ScanNet consume and label dataset-provided 3D point clouds, obtained through post processing of sensed multiview RGB-D images. They are typically trained in-domain, forego large-scale 2D pre-training and outperform alternatives that featurize the posed RGB-D multiview images instead. The gap in performance between methods that consume posed images versus post-processed 3D point clouds has fueled the belief that 2D and 3D perception require distinct model architectures. In this paper, we challenge this view and propose ODIN (Omni-Dimensional INstance segmentation), a model that can segment and label both 2D RGB images and 3D point clouds, using a transformer architecture that alternates between 2D within-view and 3D cross-view information fusion. Our model differentiates 2D and 3D feature operations through the positional encodings of the tokens involved, which capture pixel coordinates for 2D patch tokens and 3D coordinates for 3D feature tokens. ODIN achieves state-of-the-art performance on ScanNet200, Matterport3D and AI2THOR 3D instance segmentation benchmarks, and competitive performance on ScanNet, S3DIS and COCO. It outperforms all previous works by a wide margin when the sensed 3D point cloud is used in place of the point cloud sampled from 3D mesh. When used as the 3D perception engine in an instructable embodied agent architecture, it sets a new state-of-the-art on the TEACh action-from-dialogue benchmark. Our code and checkpoints can be found at the project website https://odin-seg.github.io.
@@ 1 Introduction
We observe that part of the issue lies in a key implementation detail underlying these 3D benchmark evaluations. Benchmarks like ScanNet do not actually ask methods to use RGB-D images as input, even though this is the sensor data. Instead, these benchmarks first register all RGB-D frames into a single colored point cloud and reconstruct the scene as cleanly as possible, relying on manually tuned stages for bundle adjustment, outlier rejection and meshing, and ask models to label the output reconstruction. While it is certainly viable to scan and reconstruct a room before labelling any of the objects inside, this pipeline is perhaps inconsistent with the goals of embodied vision (and typical 2D vision), which involves dealing with actual sensor data and accounting for missing or partial observations. We therefore hypothesize that method rankings will change, and the impact of 2D pre-training will become evident, if we force the 3D models to take posed RGB-D frames as input rather than pre-computed mesh reconstructions. Our revised evaluation setting also opens the door to new methods, which can train and perform inference in either single-view or multi-view settings, with either RGB or RGB-D sensors.
@@
We propose Omni-Dimensional INstance segmentation (ODIN)footnotetext: The Norse god Odin sacrificed one of his eyes for wisdom, trading one mode of perception for a more important one. Our approach sacrifices perception on post-processed meshes for perception on posed RGB-D images. , a model for 2D and 3D object segmentation and labelling that can parse single-view RGB images and/or multiview posed RGB-D images. As shown in Fig. 1, ODIN alternates between 2D and 3D stages in its architecture, fusing information in 2D within each image view, and in 3D across posed image views. At each 2D-to-3D transition, it unprojects 2D tokens to their 3D locations using the depth maps and camera parameters, and at each 3D-to-2D transition, it projects 3D tokens back to their image locations. Our model differentiates between 2D and 3D features through the positional encodings of the tokens involved, which capture pixel coordinates for 2D patch tokens and 3D coordinates for 3D feature tokens. When dealing with 2D single-view input, our architecture simply skips the 3D layers and makes a forward pass with 2D layers alone.
@@
We test ODIN in 2D and 3D instance segmentation and 3D semantic segmentation on the 2D COCO object segmentation benchmark and the 3D benchmarks of ScanNet [6], ScanNet200 [41], Matterport3D [2], S3DIS [1] and AI2THOR [25, 7]. When compared to methods using pre-computed mesh point cloud as input, our approach performs slightly worse than state-of-the-art on ScanNet and S3DIS, but better on ScanNet200 and Matterport3D. When using real sensor data as input with poses obtained from bundle reconstruction for all methods, our method performs even better, outperforming all prior work by a wide margin, in all datasets. We demonstrate that our model’s ability to jointly train on 3D and 2D datasets results in performance increase on 3D benchmarks, and also yields competitive segmentation accuracy on the 2D COCO benchmark. Our ablations show that interleaving 2D and 3D fusion operations outperforms designs where we first process in 2D and then move to 3D, or simply paint 3D points with 2D features. Stepping toward our broader goal of embodied vision, we also deploy ODIN as the 3D object segmentor of a SOTA embodied agent model [42] on the simulation benchmark TEACh [36] in the setup with access to RGB-D and pose information from the simulator, and demonstrate that our model sets a new state-of-the-art. We make our code publicly available at https://odin-seg.github.io.
@@ 2 Related Work
Early methods in 3D instance segmentation [22, 15, 3, 30, 49, 58] group their semantic segmentation outputs into individual instances. Recently, Mask2Former [4] achieved state-of-the-art in 2D instance segmentation by instantiating object queries, each directly predicting an instance segmentation mask by doing dot-product with the feature map of the input image. Inspired by it, Mask3D [44] abandons the grouping strategy of prior 3D models to use the simple decoder head of Mask2Former. MAFT [28] and QueryFormer [34] improve over Mask3D by incorporating better query initialization strategies and/or relative positional embeddings. While this shift to Mask2Former-like architecture brought the 3D instance segmentation architectures closer to their 2D counterparts, the inputs and backbones remain very different: 2D models use pre-trained backbones [16, 33], while 3D methods [44] operate over point clouds and use sparse convolution-based backbones [5], trained from scratch on small-scale 3D datasets. In this work, we propose to directly use RGB-D input and design architectures that can leverage strong 2D backbones to achieve strong performance on 3D benchmarks.
@@
Recent advancements in 2D foundation models [39, 24] have spurred efforts to apply them to 3D tasks such as point cloud classification [38, 56, 52], zero-shot 3D semantic segmentation [37, 21, 14] and more recently, zero-shot instance segmentation [47]. Commonly, these methods leverage 2D foundation models to featurize RGB images, project 3D point clouds onto these images, employ occlusion reasoning using depth and integrate features from all views through simple techniques like mean-pooling. Notably, these approaches predominantly focus on semantic segmentation, emphasizing pixel-wise labeling, rather than instance labeling, which necessitates cross-view reasoning to associate the same object instance across multiple views. OpenMask3D [47] is the only method that we are aware of that attempts 3D instance segmentation using 2D foundation models, by training a class-agnostic 3D object segmentor on 3D point clouds and labelling it utilizing CLIP features. Despite their effectiveness in a zero-shot setting, they generally lag behind SOTA 3D supervised methods by 15-20%. Rather than relying on features from foundation models, certain works [12, 10] create 3D pseudo-labels using pre-trained 2D models. Another line of work involves fitting Neural-Radiance Fields (NeRFs), incorporating features from CLIP [23, 48] or per-view instance segmentations from state-of-the-art 2D segmentors [46]. These approaches require expensive per-scene optimization that prohibits testing on all test scenes to compare against SOTA 3D discriminative models. Instead of repurposing 2D foundation models for 3D tasks, Omnivore [13] proposes to build a unified architecture that can handle multiple visual modalities like images, videos and single-view RGB-D image but they only show results for classification tasks. We similarly propose a single unified model capable of performing both single-view 2D and multi-view 3D instance and semantic segmentation tasks while utilizing pre-trained weights for the majority of our architecture.
@@ 3 Method
$\textrm{query}_{\textrm{pos}}$ $=\textrm{MLP}(0);$ (1)
$\textrm{key}_{\textrm{pos}}$ $=\textrm{MLP}(p_{i}-p_{j}),$ (2)
@@
Open vocabulary class decoder: Drawing inspiration from prior open-vocabulary detection methods [19, 29, 61], we introduce an alternative classification head capable of handling an arbitrary number of semantic classes. This modification is essential for joint training on multiple datasets. Similar to BUTD-DETR [19] and GLIP [29], we supply the model with a detection prompt formed by concatenating object categories into a sentence (e.g., “Chair. Table. Sofa.") and encode it using RoBERTa [32]. In the query-refinement block, queries additionally attend to these text tokens before attending to the upsampled feature maps. For semantic class prediction, we first perform a dot-product operation between queries and language tokens, generating one logit per token in the detection prompt. The logits corresponding to prompt tokens for a specific object class are then averaged to derive per-class logits. This can handle multi-word noun phrases such as “shower curtain", where we average the logits corresponding to “shower" and “curtain". The segmentation masks are predicted by a pixel-/point-wise dot-product, in the same fashion as described earlier.
@@ 4.1 Evaluation on 3D benchmarks
Table 1: Evaluation on 3D Benchmarks (§ = trained by us using official codebase).
(a) ScanNet Instance Segmentation Task.
@@
Sensor RGBD Point Cloud Mask3D§ [44] 43.9 60.0 69.9
ODIN-ResNet50 (Ours) 47.8 69.8 83.6
ODIN-Swin-B (Ours) 50.0 71.0 83.6
@@
(b) ScanNet Semantic Segmentation Task.
@@
ODIN-ResNet50 (Ours) 73.3
ODIN-Swin-B (Ours) 77.8
@@
(c) ScanNet200 Instance Segmentation Task.
@@
Sensor RGBD Point Cloud Mask3D [44] § 15.5 21.4 24.3
ODIN-ResNet50 (Ours) 25.6 36.9 43.8
ODIN-Swin-B (Ours) 31.5 45.3 53.1
@@
(d) ScanNet200 Semantic Segmentation Task.
@@
Sensor RGBD Point Cloud ODIN-ResNet50 (Ours) 35.8
ODIN-Swin-B (Ours) 40.5
@@
Datasets: First, we test our model on 3D instance and semantic segmentation in the ScanNet [6] and ScanNet200 [41] benchmarks. The goal of these benchmarks is to label the point cloud extracted from the 3D mesh of a scene reconstructed from raw sensor data. ScanNet evaluates on 20 common semantic classes, while ScanNet200 uses 200 classes, which is more representative of the long-tailed object distribution encountered in the real world. We report results on the official validation split of these datasets here and on the official test split in the supplementary.
@@
ODIN sets a new SOTA in semantic segmentation on ScanNet (Tab. 1(b)) outperforming all methods on all setups including the models trained on the sensor, rendered and mesh sampled point clouds.
ODIN sets a new instance segmentation SOTA on the long-tailed ScanNet200 dataset (Tab. 1(c)) outperforming SOTA 3D models on all setups including the models trained on mesh-sampled point cloud, especially by a large margin in mAP25 metric, while exclusively utilizing sensor RGB-D data. This highlights the contribution of 2D features, particularly in detecting a long tail of class distribution where limited 3D data is available. We show more detailed results with performance on the head, common and tail classes in the appendix ( Sec. A.3).
@@
Table 2: AI2THOR Semantic and Instance Segmentation.
@@
ODIN-ResNet50 (Ours) 63.8 73.8 80.2 71.5
ODIN-Swin-B (Ours) 64.3 73.7 78.6 71.4
@@ 4.2 Evaluation on multiview RGB-D in simulation
Using the AI2THOR [25] simulation environment with procedural homes from ProcThor [7], we collected RGB-D data for 1500 scenes (1200 training, 300 test) of similar size as ScanNet (more details in appendix,  Appendix B). We train and evaluate our model and SOTA Mask3D [44] on the unprojected RGB-D images. As shown in Tab. 2, our model outperforms Mask3D by 3.7% mAP, showing strong performance in a directly comparable RGB-D setup. It suggests that current real-world benchmarks may restrain models that featurizes RGB-D sensor point clouds due to misalignments. We hope this encourages the community to also focus on directly collecting, labeling, and benchmarking RGB-D sensor data.
@@ 4.3 Embodied Instruction Following
Mask3D [44] 43.9 60.0 69.9
Mask2Former [4] 43.7
ODIN (trained in 2D) 43.6
ODIN (trained in 3D) 47.8 69.8 83.6
ODIN (trained jointly) 49.1 70.1 83.1 41.2
@@
Model mAP mIoU
ODIN (Ours) 47.8 73.3
No 3D Fusion 39.3 73.2
No interleaving 41.7 73.6
@@
Model mAP mIoU
ODIN (Ours) 47.8 73.3
Only pre-trained backbone 42.3 72.9
No pre-trained features 41.5 68.6
@@
Model ResNet50 Swin-B
@@
ODIN (Ours) 47.8 73.3 50.0 77.8
With frozen backbone 46.7 74.3 46.2 75.9
We apply ODIN in the embodied setups of TEACh [36] and ALFRED [45] where agents have access to RGB, depth and camera poses and need to infer and execute task and action plans from dialogue segments and instructions, respectively. These agents operate in dynamic home environments and cannot afford expensive mesh building steps. Detecting objects well is critical for task success in both cases. Prior SOTA methods [42, 36] run per-view 2D instance segmentation models [4, 9] and link the detected instances using simple temporal reasoning regarding spatial and appearance proximity. Instead, ODIN processes its last $N$ egocentric views and segments objects instances directly in 3D. We equip HELPER [42], a state-of-the-art embodied model, with ODIN as its 3D object detection engine. We evaluate using Task Sucess Rate (SR) which checks if the entire task is executed successfully, and Goal Conditioned Success Rate (GC) which checks the proportion of satisfied subgoals across all episodes [45, 36]. We perform evaluation on "valid-seen" (houses similar to the training set) and "valid-unseen" (dissimilar houses) splits. In Tab. 3, we observe that HELPER with ODIN as its 3D object detector significantly outperforms HELPER that uses the original 2D detection plus linking perception pipeline.